* Remote participants by eZuce: None
Report of work (*:responsible person)
- ETMX had a jump at 5/29 and stuck. There was an earthquake on the same day but the time was not matched by ~hour.
- MCO TFs were measured, and behaviors did not change.
- TFs of P to P and Y to Y look fine.
- MC alignment is still not well. Beam will be traced with higher power.
Work plan (*:responsible person)
- Initial alignment for PR2 will be discussed.
- ETMX had a jump at 5/29 and stuck.
- MCO had a big coupling between L and P,Y.